OpenStudy (anonymous):

2. A prism has total original surface area of 198m2 and volume of 162 m3. a) What is the surface area to volume ratio? (reduce) b) If the length, width, and height are reduced to one third of their original sizes, what is the multiplier? c) What is the new surface area after the reduction? (show your work!) d) What is the new volume after the reduction? (show your work!) e) If the length, width, and height of the original are increased to twice their original sizes, what will be the multiplier? f) If the length, width, and height of the original are increased to tw

OpenStudy (anonymous):

a. Ratio: Surface/volume = 198/162 = 33/17 b. The multiplier is: 1/3 for size; 1/9 for surfaces; and 1/27 for volumes. c. Surface after reduction: 198 x (1/9) = 22 d. Volume after reduction: 162 *(1/27) = 6 e. Multiplier: for size : 1/2; for surface : 1/4 ; for volume 1/8